Solve the problem by using a system of equations:
The sum of 2 numbers is 188 and their difference is 2. Find the numbers.

Answer by stanbon(75887)

The sum of 2 numbers is 188 and their difference is 2.
x+y=188
x-y=2
--------
Add the equations to get:
2x=190
x=95
Substitute to solve for y:
95+y=188
y= 93
========
Cheers,
Stan H.